1. A wheel with a diameter of 80cm speeds up from 30 rad/s to 80 rad/s. The linear acceleration at the edge of the wheel is 15 m/s2. How many revolutions will the wheel go through during this period?
1
Expert's answer
2021-11-24T11:48:37-0500

Let's first find the angular acceleration of the wheel:


"a=r\alpha,""\alpha=\dfrac{a}{r}=\dfrac{15\ \dfrac{m}{s^2}}{0.4\ m}=37.5\ \dfrac{rad}{s^2}."


Also, we need to find the time that the wheel takes to accelerate:


"t=\dfrac{\omega_f-\omega_i}{\alpha}=\dfrac{80\ \dfrac{rad}{s}-30\ \dfrac{rad}{s}}{37.5\ \dfrac{rad}{s^2}}=1.33\ s."

Then, we can find the angular displacement of the wheel:


"\theta=\omega_it+\dfrac{1}{2}\alpha t^2,""\theta=30\ \dfrac{rad}{s}\times1.33\ s+\dfrac{1}{2}\times37.5\ \dfrac{rad}{s^2}\times(1.33\ s)^2=73\ rad."

Finally, we can find the number of revolutions that the wheell will go through during  this period:


"n=\dfrac{\theta}{2\pi\ \dfrac{rad}{rev}}=\dfrac{73\ rad}{2\pi\ \dfrac{rad}{rev}}=12\ rev."